OpenStudy (anonymous):

FAN AND MEDAL!!! Solve for x: -2(x + 3) = -2(x + 1) - 4 I need the work as well!

OpenStudy (anonymous):

it is an identity. all real numbers are solutions.

OpenStudy (anonymous):

How do I figure that out, though?

OpenStudy (anonymous):

@PinkSapphire Do you know how to do this?

OpenStudy (pinksapphire):

-2(x+3)=-2(x+1)-4 _2x-6=-2x-2-4 Distribute first. _2x-6=-2x-2-4 Add 2x to both sides and they cancel each other out. +2x +2x -6=-2-4 -6=-6 Add 6 to both sides. +6 +6 ?=? All numbers are solutions.

OpenStudy (anonymous):

thts wht i did.i was posting my work but labtop is 2 slow

OpenStudy (anonymous):

Okay. Thank You BOTH! You are lifesavers :)

OpenStudy (anonymous):

no prob.